OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8).

OpenStudy (anonymous):

Use distance formula Sqrt[(x-x1)^2+(y-y1)^2]

OpenStudy (anonymous):

\[\sqrt{(-7+2)^{2} + (-8+3)^{2}} = \sqrt{25+25} = \sqrt{50}\]